# Spectrum of slip dynamics, scaling & statistical laws emerge from simplified model of fault and damage zone architecture

### v1

@@ Acknowledgements
MA, NK, CV, JC, AG and HSB gratefully acknowledge the European Research Council (ERC) for its full support of this work through the PERSISMO grant (No. 865411). MYT acknowledges support from the Agence National de la Recherche (ANR) IDEAS contract ANR-19-CE31-0004-01. PR acknowledges support from the GPX program,funded by the Agence National de la Recherche (ANR), CGG, TOTAL, and Schlumberger for his PhD fellowship and the European Research Council (ERC) Starting Grant 101040600 (HYQUAKE). NK acknowledges funding from Horizon Europe (ChEESE-2P project, grant agreement No. 101093038) for partial support in working on this manuscript. This paper has benefited from kind discussions and feedbacks from Satoshi Ide, Hideo Aochi, Romain Jolivet and Raul Madariaga. The numerical simulations presented in this study were performed on the MADARIAGA cluster, also supported by the ERC PERSISMO grant. We also used LLM models like GPT-4 and Claude to debug and optimize our codes.
@@
The datasets generated and analyzed during this study, and codes used to analyze, are available on the Zenodo repository: https://doi.org/10.5281/zenodo.17061565.
Appendix A Model of earthquake cycle
A1 Governing equation
@@ Appendix A Model of earthquake cycle
$\tau^{f}(s)=\tau_{t}^{\text{el}}(s)+\tau^{\text{rad}}(s)+\tau^{\text{load}}(s)$ (A1)
@@
$\tau^{f}(s,t)=-f(V,\theta)\sigma_{n}(s,t)=-a\sigma_{n}(s,t)\sinh^{-1}\left[\frac{V(s,t)}{2V_{0}}\exp\left\{\frac{f_{0}+b\ln[V_{0}\theta(s,t)/d_{c}]}{a}\right\}\right]$ (A2)
@@
$\dfrac{d\theta(s,t)}{dt}=1-\frac{\theta(s,t)V(s,t)}{d_{c}}$ (A3)
$a$ is the direct effect parameter that governs the instantaneous change in friction with a change in slip rate, $b$ is the the evolution effect parameter which controls how friction evolves over time via changes in the state variable, $d_{c}$ is the characteristic slip for state evolution, $V_{0}$ is reference slip rate and $f_{0}$ is the reference friction at $_{0}. VV$ and $\theta$ represent respectively the slip velocity and the state variable. $\sigma_{n}$ is the normal traction on the fault and $d_{c}$ is the critical slip distance. The term $f_{0}+b\log(V_{0}\theta/d_{c})$ is sometimes denoted by $\Psi$.
A3 Elastostatic response of the fault due to a slip distribution
@@
$\tau_{t}^{\text{el}}(s)=\mathbf{t}(s)\cdot\overline{\overline{\sigma}}(\mathbf{y}(s))\cdot\mathbf{n}(s)=\int_{\Gamma}\!K_{\text{grad}}^{t}(s,\xi)\frac{\partial}{\partial\xi}\Delta u_{t}(\xi)\mathrm{d}\xi+\int_{\Gamma}\!K_{\text{curv}}^{t}(s,\xi)\kappa^{t}(\xi)\Delta u_{t}(\xi)\mathrm{d}\xi$ (A4)
$\sigma_{n}^{\text{el}}(s)=\mathbf{n}(s)\cdot\overline{\overline{\sigma}}(\mathbf{y}(s))\cdot\mathbf{n}(s)=\int_{\Gamma}\!K_{\text{grad}}^{n}(s,\xi)\frac{\partial}{\partial\xi}\Delta u_{t}(\xi)\mathrm{d}\xi+\int_{\Gamma}\!K_{\text{curv}}^{n}(s,\xi)\kappa^{t}(\xi)\Delta u_{t}(\xi)\mathrm{d}\xi$ (A5)
In order to set the model of earthquake cycle, we chose the compression negative sign convention. $s$ is the curvilinear location along the fault, $\Gamma\equiv\mathbf{y}(s)$, at which the stress is evaluated. $\xi$ is a curvilinear location, at which a source of stress is located. $\mathbf{n}(\xi)$ and $\mathbf{t}(\xi)$ are respectively the normal and tangential vector to the fault at point $\xi$. Their component in the global coordinate system are written with the corresponding subscript. Finally, $K^{t}_{\text{grad}}$ is the kernel the kernel for the tangential traction, and $K^{n}_{\text{grad}}$ and $K^{n}_{\text{curv}}$ are the kernel for the normal traction. $K^{t}_{\text{grad}}$ and $K^{t}_{\text{curv}}$ are the kernel for the tangential traction associated respectively to the effect of the gradient of tangential slip and local curvature that multiplies the tangential slip. They both derive from the fact that a derivative of the tangential slip vector along the fault is:
$\begin{split}\frac{\partial}{\partial\xi}[\Delta u_{t}(\xi)\mathbf{t}(\xi)]&=\frac{\partial}{\partial\xi}[\Delta u_{t}(\xi)]\mathbf{t}(\xi)+\Delta u_{t}(\xi)\frac{\partial}{\partial\xi}\mathbf{t}(\xi)\\ &=\frac{\partial}{\partial\xi}[\Delta u_{t}(\xi)]\mathbf{t}(\xi)+\Delta u_{t}(\xi)\kappa^{t}(\xi)\mathbf{n}(\xi)\end{split}$ (A6)
@@
Similarly, $K^{n}_{\text{grad}}$ and $K^{n}_{\text{curv}}$ are the kernel for the normal traction. Mode details about the derivation of these equations can be found in [161]. It should be noted here that these kernels take into account the full elastic interaction between all faults in the system, including the main fault and all off-fault fractures.
A4 Radiation damping term
@@
$\tau^{\text{rad}}(s)=-\dfrac{\mu}{2c_{s}}V(s)$ (A7)
We use this term together with the static kernel to account for some dynamics in the system. Without this term, the slip on the fault during one event would be unbounded [154].
A5 Set of ordinary differential equations for Quasi-Dynamic earthquake cycle models
Balance of forces requires the strength of the fault to be equal to the elastic shear traction (due to slip) plus the far-field loading traction plus radiation damping term (eq. A1).
By differentiating eq. A1 with time $t$, it can be recast into a set of coupled ODEs. The slip acceleration becomes:
$\dfrac{dV(s,t)}{dt}=\frac{\dot{\tau}^{\text{load}}(s,t)+\dot{\tau}_{t}^{\text{el}}(s,t)+\dfrac{\partial f(V,\theta)}{\partial\theta}\dfrac{d\theta}{dt}\sigma_{n}(s,t)+f(V,\theta)\dot{\sigma}_{n}(s,t)}{\dfrac{\mu}{2c_{s}}-\dfrac{\partial f(V,\theta)}{\partial V}\sigma_{n}(s,t)}$ (A8)
@@
$\dot{\tau}_{t}^{\text{el}}(s,t)=\int_{\Gamma}\!K_{\text{grad}}^{t}(s,\xi)\frac{\partial}{\partial\xi}V(\xi,t)\mathrm{d}\xi+\int_{\Gamma}\!K_{\text{curv}}^{t}(s,\xi)\kappa^{t}(\xi)V(\xi,t)\mathrm{d}\xi$ (A9)
@@
$\dfrac{d\sigma_{n}(s,t)}{dt}=\int_{faults}K_{\text{grad}}^{\text{n}}(s,\xi)\dfrac{\partial}{\partial\xi}V(\xi,t)\mathrm{d}\xi+\int_{faults}K_{\text{curv}}^{\text{n}}(s,\xi)\kappa^{\text{t}}(\xi)V(\xi,t)\mathrm{d}\xi+\dot{\sigma}_{n}^{load}(s,t)$ (A10)
And we recall the evolution of the state variable eq. A3. This set of ODEs is then solved at each centre of element, using the Runge-Kutta 45 adaptive time step ODE solver algorithm [37].
A6 Numerical discretisation of the Boundary Integral Equation
@@
$\Delta u(s)=\sum_{j=1}^{N}\Delta u(s_{j})[\mathcal{H}(s-s_{j}+\Delta s/2)-\mathcal{H}(s-s_{j}-\Delta s/2)]$ (A11)
@@
$\mathbf{t}(s)=\sum_{j=1}^{N}\mathbf{t}(s_{j})[\mathcal{H}(s-s_{j}+\Delta s/2)-\mathcal{H}(s-s_{j}-\Delta s/2)]$ (A12)
Then the boundary integral equations A4 and A5 become a summation:
@@
A7 Hierarchical Matrices
Hierarchical matrices have already been used in the context of quasi dynamic modeling of faults [31, 159, 42]. Hierarchical matrices provide data-sparse approximations of non-sparse (dense) matrices [74, 29]. Hierarchical matrices provide an approximation requiring only $\mathcal{O}(N\log{N})$ units of storage (instead of $\mathcal{O}(N^{2})$). The two advantages of using H-matrices are that it reduces the memory for saving the full matrix of interaction, and it is also reducing the number of operations to perform the matrix-vector products. The use of H matrices is extremely pertinent in quasidynamic models. The kernel matrix of stress interaction A6 is not changing over time, making it a suitable matrix for compression, since the most expensive part, the construction of the data-sparse representation, is performed only once, while subsequent matrix-vector products are extremely fast. The accuracy of this method is completely controlled by the tolerance set for the compression, particularly if the required minimum accuracy approaches machine precision. The mathematical formulation of hierarchical matrices is beyond the scope of this paper but is very well explained in [29, 52, 159, 42].
A8 Frictional length scales
@@
$k_{c}=\sigma_{n}\frac{b-a}{d_{c}}=\frac{\mu}{L_{nuc}}=k_{fault}$ (A14)
@@
$\begin{cases}L_{nuc}=2\times 1.3774L_{b}&0\leq a/b<0.3781\\ L_{nuc}=2\times\dfrac{L_{b}}{\pi(1-a/b)^{2}}&a/b\to 1\end{cases}$ (A17)
@@
Appendix B Catalog building algorithm
@@
Figure B1: Steps involved in catalog generation from slip rate matrix. (a) Slip rate matrix (time, space). (b) Binary image after thresholding slip rates with Labeled connected components. (c) Bounding boxes around each connected component, representing individual events.

### v2

@@ Acknowledgements
MA, NK, CV, JC, AG and HSB gratefully acknowledge the European Research Council (ERC) for its full support of this work through the PERSISMO grant (No. 865411). MYT acknowledges support from the Agence National de la Recherche (ANR) IDEAS contract ANR-19-CE31-0004-01. PR acknowledges support from the GPX program,funded by the Agence National de la Recherche (ANR), CGG, TOTAL, and Schlumberger for his PhD fellowship and the European Research Council (ERC) Starting Grant 101040600 (HYQUAKE). NK acknowledges funding from Horizon Europe (ChEESE-2P project, grant agreement No. 101093038) for partial support in working on this manuscript. This paper has benefited from kind discussions and feedbacks from Satoshi Ide, Hideo Aochi, Romain Jolivet and Raul Madariaga. The authors also acknowledge the feedback from the editor R. Abercrombie, the associate editor A-A. Gabriel, Prof. Y. Huang and an anonymous reviewer. The numerical simulations presented in this study were performed on the MADARIAGA cluster, also supported by the ERC PERSISMO grant. We also used LLM models like GPT-4 and Claude to debug and optimize our codes.
@@
The datasets generated and analyzed during this study, and codes used to analyze, are available on the Zenodo repository: https://doi.org/10.5281/zenodo.18599125.
Supplementary Information for
“Fault volume digital twin to reproduce the full slip spectrum, scaling and statistical laws”
M. Almakari1,†, N. Kheirdast1,†, C. Villafuerte1,∗, M. Y. Thomas2, P. Dubernet1, J. Cheng1,§, A. Gupta1, P. Romanet3,4, S. Chaillat5, H. S. Bhat1,¶
  1. 1.
    Laboratoire de Géologie, Ecole Normale Superieure, CNRS-UMR 8538, PSL Research University, Paris, France
  2. 2.
    Université de Rennes, CNRS, Géosciences Rennes, CNRS-UMR 6118, Rennes
  3. 3.
    Department of Earth Sciences, La Sapienza University of Rome, Piazzale Aldo Moro 5, 00185 Roma, Italy
  4. 4.
    Université Côte d‘Azur, CNRS, IRD, Observatoire de la Côte d’Azur, Géoazur, Sophia-Antipolis, 06560 Valbonne, France
  5. 5.
    Laboratoire POEMS, CNRS-INRIA-ENSTA Paris, Institut Polytechnique de Paris
footnotetext: $\dagger$ M. Almakari and N. Kheirdast contributed equally to this work. * Currently at Instituto de Geofisica, Universidad Nacional Autonoma de México. $\S$ Currently at Division of Geological and Planetary Sciences, California Institute of Technology. $\P$ Corresponding author: harshasbhat@gmail.com
Introduction
This Supplementary Information provides detailed technical documentation for the numerical methods, algorithms, and additional results supporting the main manuscript. The document is organized as follows:
Section S1 describes the quasi-dynamic earthquake cycle model
Section S2 presents the catalog building algorithm used to identify and characterize individual slip events from spatiotemporal slip rate matrices.
Supplementary Figures:
Figure S1 illustrates the catalog generation workflow.
Figure S2 demonstrates spatial segregation of fast and slow events in a fault system with off-fault fractures parallel to the main fault, using the same configuration as the main manuscript.
Figure S3 presents statistical properties and scaling laws
Figure S4 shows moment-duration scaling relationships with different detection thresholds.
S1 Model of earthquake cycle
S1.1 Governing equation
@@ Introduction
$\tau^{f}(s)=\tau_{t}^{\text{el}}(s)+\tau^{\text{rad}}(s)+\tau^{\text{load}}(s)$ (S1.1)
@@
$\tau^{f}(s,t)=-f(V,\theta)\sigma_{n}(s,t)=-a\sigma_{n}(s,t)\sinh^{-1}\left[\frac{V(s,t)}{2V_{0}}\exp\left\{\frac{f_{0}+b\ln[V_{0}\theta(s,t)/d_{c}]}{a}\right\}\right]$ (S1.2)
@@
$\dfrac{d\theta(s,t)}{dt}=1-\frac{\theta(s,t)V(s,t)}{d_{c}}$ (S1.3)
$a$ is the direct effect parameter that governs the instantaneous change in friction with a change in slip rate, $b$ is the evolution effect parameter which controls how friction evolves over time via changes in the state variable, $d_{c}$ is the characteristic slip for state evolution, $V_{0}$ is reference slip rate and $f_{0}$ is the reference friction at $_{0}. VV$ and $\theta$ represent respectively the slip velocity and the state variable. $\sigma_{n}$ is the normal traction on the fault and $d_{c}$ is the critical slip distance. The term $f_{0}+b\log(V_{0}\theta/d_{c})$ is sometimes denoted by $\Psi$.
S1.3 Elastostatic response of the fault due to a slip distribution
@@
$\tau_{t}^{\text{el}}(s)=\mathbf{t}(s)\cdot\overline{\overline{\sigma}}(\mathbf{y}(s))\cdot\mathbf{n}(s)=\int_{\Gamma}\!K_{\text{grad}}^{t}(s,\xi)\frac{\partial}{\partial\xi}\Delta u_{t}(\xi)\mathrm{d}\xi+\int_{\Gamma}\!K_{\text{curv}}^{t}(s,\xi)\kappa^{t}(\xi)\Delta u_{t}(\xi)\mathrm{d}\xi$ (S1.4)
$\sigma_{n}^{\text{el}}(s)=\mathbf{n}(s)\cdot\overline{\overline{\sigma}}(\mathbf{y}(s))\cdot\mathbf{n}(s)=\int_{\Gamma}\!K_{\text{grad}}^{n}(s,\xi)\frac{\partial}{\partial\xi}\Delta u_{t}(\xi)\mathrm{d}\xi+\int_{\Gamma}\!K_{\text{curv}}^{n}(s,\xi)\kappa^{t}(\xi)\Delta u_{t}(\xi)\mathrm{d}\xi$ (S1.5)
In order to set the model of earthquake cycle, we chose the compression negative sign convention. $s$ is the curvilinear location along the fault, $\Gamma\equiv\mathbf{y}(s)$, at which the stress is evaluated. $\xi$ is a curvilinear location, at which a source of stress is located. $\mathbf{n}(\xi)$ and $\mathbf{t}(\xi)$ are respectively the normal and tangential vector to the fault at point $\xi$. Their component in the global coordinate system are written with the corresponding subscript. Finally, $K^{t}_{\text{grad}}$ is the kernel for the tangential traction, and $K^{n}_{\text{grad}}$ and $K^{n}_{\text{curv}}$ are the kernel for the normal traction. $K^{t}_{\text{grad}}$ and $K^{t}_{\text{curv}}$ are the kernels for the tangential traction associated respectively to the effect of the gradient of tangential slip and local curvature that multiplies the tangential slip. They both derive from the fact that a derivative of the tangential slip vector along the fault is:
$\begin{split}\frac{\partial}{\partial\xi}[\Delta u_{t}(\xi)\mathbf{t}(\xi)]&=\frac{\partial}{\partial\xi}[\Delta u_{t}(\xi)]\mathbf{t}(\xi)+\Delta u_{t}(\xi)\frac{\partial}{\partial\xi}\mathbf{t}(\xi)\\ &=\frac{\partial}{\partial\xi}[\Delta u_{t}(\xi)]\mathbf{t}(\xi)+\Delta u_{t}(\xi)\kappa^{t}(\xi)\mathbf{n}(\xi)\end{split}$ (S1.6)
@@
Similarly, $K^{n}_{\text{grad}}$ and $K^{n}_{\text{curv}}$ are the kernels for the normal traction. Mode details about the derivation of these equations can be found in [150]. It should be noted here that these kernels take into account the full elastic interaction between all faults in the system, including the main fault and all off-fault fractures.
S1.4 Radiation damping term
@@
$\tau^{\text{rad}}(s)=-\dfrac{\mu}{2c_{s}}V(s)$ (S1.7)
where $c_{s}$ is the shear wave speed. We use this term together with the static kernel to account for some dynamics in the system. Without this term, the slip on the fault during one event would be unbounded [144].
S1.5 Set of ordinary differential equations for Quasi-Dynamic earthquake cycle models
Balance of forces requires the strength of the fault to be equal to the elastic shear traction (due to slip) plus the far-field loading traction plus radiation damping term (eq. S1.1).
By differentiating eq. S1.1 with time $t$, it can be recast into a set of coupled ODEs. The slip acceleration becomes:
$\dfrac{dV(s,t)}{dt}=\frac{\dot{\tau}^{\text{load}}(s,t)+\dot{\tau}_{t}^{\text{el}}(s,t)+\dfrac{\partial f(V,\theta)}{\partial\theta}\dfrac{d\theta}{dt}\sigma_{n}(s,t)+f(V,\theta)\dot{\sigma}_{n}(s,t)}{\dfrac{\mu}{2c_{s}}-\dfrac{\partial f(V,\theta)}{\partial V}\sigma_{n}(s,t)}$ (S1.8)
@@
$\dot{\tau}_{t}^{\text{el}}(s,t)=\int_{\Gamma}\!K_{\text{grad}}^{t}(s,\xi)\frac{\partial}{\partial\xi}V(\xi,t)\mathrm{d}\xi+\int_{\Gamma}\!K_{\text{curv}}^{t}(s,\xi)\kappa^{t}(\xi)V(\xi,t)\mathrm{d}\xi$ (S1.9)
@@
$\dfrac{d\sigma_{n}(s,t)}{dt}=\int_{faults}K_{\text{grad}}^{\text{n}}(s,\xi)\dfrac{\partial}{\partial\xi}V(\xi,t)\mathrm{d}\xi+\int_{faults}K_{\text{curv}}^{\text{n}}(s,\xi)\kappa^{\text{t}}(\xi)V(\xi,t)\mathrm{d}\xi+\dot{\sigma}_{n}^{load}(s,t)$ (S1.10)
And we recall the evolution of the state variable eq. S1.3. This set of ODEs is then solved at each centre of element, using the Runge-Kutta 45 adaptive time step ODE solver algorithm [33].
S1.6 Numerical discretisation of the Boundary Integral Equation
@@
$\Delta u(s)=\sum_{j=1}^{N}\Delta u(s_{j})[\mathcal{H}(s-s_{j}+\Delta s/2)-\mathcal{H}(s-s_{j}-\Delta s/2)]$ (S1.11)
@@
$\mathbf{t}(s)=\sum_{j=1}^{N}\mathbf{t}(s_{j})[\mathcal{H}(s-s_{j}+\Delta s/2)-\mathcal{H}(s-s_{j}-\Delta s/2)]$ (S1.12)
Then the boundary integral equations S1.4 and S1.5 become a summation:
@@
S1.7 Hierarchical Matrices
Hierarchical matrices have already been used in the context of quasi dynamic modeling of faults [27, 152, 38]. Hierarchical matrices provide data-sparse approximations of non-sparse (dense) matrices [65, 25]. Hierarchical matrices provide an approximation requiring only $\mathcal{O}(N\log{N})$ units of storage (instead of $\mathcal{O}(N^{2})$). The two advantages of using H-matrices are that it reduces the memory for saving the full matrix of interaction, and it is also reducing the number of operations to perform the matrix-vector products. The use of H matrices is extremely pertinent in quasidynamic models. The kernel matrix of stress interaction S1.6 is not changing over time, making it a suitable matrix for compression, since the most expensive part, the construction of the data-sparse representation, is performed only once, while subsequent matrix-vector products are extremely fast. The accuracy of this method is completely controlled by the tolerance set for the compression, particularly if the required minimum accuracy approaches machine precision. The mathematical formulation of hierarchical matrices is beyond the scope of this paper but is very well explained in [25, 46, 152, 38].
S1.8 Frictional length scales
@@
$k_{c}=\sigma_{n}\frac{b-a}{d_{c}}=\frac{\mu}{L_{nuc}}=k_{fault}$ (S1.14)
@@
$\begin{cases}L_{nuc}=2\times 1.3774L_{b}&0\leq a/b<0.3781\\ L_{nuc}=2\times\dfrac{L_{b}}{\pi(1-a/b)^{2}}&a/b\to 1\end{cases}$ (S1.17)
@@
S2 Catalog building algorithm
@@
Figure S1: Steps involved in catalog generation from slip rate matrix. (a) Slip rate matrix (time, space). (b) Binary image after thresholding slip rates with Labeled connected components. (c) Bounding boxes around each connected component, representing individual events.
@@
Figure S2: Segregation of slow and fast events at two separated regions of the fault. The geometrical configuration is the same as in Figure 1a except that the off-fault fractures are parallel to the main fault.
Figure S3: (a) Omori Law: Seismicity rate decreases over time following the mainshock. Each gray curve represents one earthquake cycle with one mainshock on the main fault, the blue curve represents the stacked sequences of all the seismic cycles in this simulation case (b) Magnitude-frequency distribution of the cataloged fast ruptures follows the Gutenberg-Richter distribution. Red and blue colors represent fast and slow ruptures respectively, while black color represents the total catalog (c) Inverse Omori: Seismicity rate increases inversely with time as the main rupture approaches. Each gray curve represents one earthquake cycle with one mainshock on the main fault, the blue curve represents the stacked sequences of all the seismic cycles in this simulation case (d) Scaling laws of inferred seismicity: Moment area scaling. Events are color-coded based on rupture velocity $v_{r}$.
Figure S4: Scaling laws of inferred seismicity: Moment-duration scaling with a threshold of detection of slow ruptures at (a) $10^{-6}$ m/s and (b) $10^{-8}$ m/s. Events are color-coded based on rupture velocity $v_{r}$.